Circular No. 8090

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003br IN MCG –05-34-18 Further to IAUC 8089, B. Swift and W. Li report the LOTOSS dis-

covery of an apparent supernova (mag ∼ 17.4) on unfiltered KAIT images taken on Mar. 7.5 and 8.5 UT. SN 2003br is located at α = 14^{h}34^{m}34^{s}.98, δ = −27^{o}59′48′′.8 (equinox 2000.0), which is 2′′.3 west and 5′′.5 south of the nucleus of MCG –05-34-18. A KAIT image taken on 2002 July 1.3 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2003bs Further to IAUC 8089, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of a supernova (mag 17.7) on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope on Mar. 3.48, 3.49, and 3.50 UT. The new object is located at α = 10^{h}29^{m}44^{s}.12, δ = +25^{o}37′11′′.6 (equinox 2000.0), which is 0′′.8 west and 0′′.6 north of the center of the apparent host galaxy. No point source is visible at this location on Palomar images from Jan. 28.45–28.49 (limiting mag 20.1). D. A. Howell, Lawrence Berkeley National Laboratory, reports that a spectrum (range 360–900 nm) of SN 2003bs, obtained on Mar. 8.32 with the Kitt Peak 2.1-m reflector (+ GCAM), shows it to be a type-Ia supernova, ≈ 1 week past peak brightness. The host galaxy is very faint, and the redshift (z = 0.05) was determined by comparison of the spectral features of the supernova to the strongly resembling features of SN 2001ay (cf. IAUC 7612).

COMET C/2002 X5 (KUDO-FUJIKAWA) Ephemeris extension to IAUC 8046 (from orbital elements on MPEC

2003-E20; magnitude parameters H = 8.0, n = 3):

2003 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 12 3^{h}58.^{m}31 −24^{o}29.^{′}1 1.082 1.169 68.^{o}4 52.^{o}2  8.7
17 4 23.89 −19 52.3 1.191 1.270 70.4 47.5  9.2
22 4 44.47 −15 55.2 1.313 1.368 71.2 43.6  9.6
27 5 01.55 −12 34.9 1.442 1.463 71.0 40.2 10.0
Apr. 1 5 16.12 − 9 46.3 1.577 1.556 70.2 37.2 10.4
6 5 28.85 − 7 24.2 1.716 1.646 68.9 34.5 10.8
11 5 40.22 − 5 24.1 1.857 1.734 67.2 32.2 11.1
16 5 50.53 − 3 42.1 1.999 1.820 65.1 30.0 11.5
21 6 00.01 − 2 15.2 2.141 1.905 62.8 28.0 11.8
2003 March 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green